OpenStudy (anonymous):

solve the system algebraically y=x^2, 9x-y=20

OpenStudy (anonymous):

second equation says \[y=9x-20\] replace in first equation to get \[9x-20=x^2\] \[x^2-9x+20=0\] \[(x-4)(x-5)=0\] \[x=4\] \[x=5\]